Suppose there are three types of Apples A, B and C being sold and consumed. The demand and supply equations for each type are:


Da=20-2Pa+4Pb+Pc; Sa=4Pa-5; Db=10+3Pa-5Pb+2Pc ;Sb=3Pb-7;Dc=70+4Pa+2Pb-5Pc ;Sc=5Pc-16


Determine equilibrium prices and quantities using Cramer’s rule.


• Calculate the elasticity of demand for B with respect to prices of variety A, B and C and interpret the economic meaning of the results.

1
Expert's answer
2021-12-17T08:41:01-0500

"D_A= 20 – 2P_A + 4P_B + P_C \\

S_A=4P_A-5 \\

D_B=10+3P_A-5P_B+2P_C \\

S_B=3P_B-7 \\

D_C=70+4P_A+2P_B-5P_C \\

S_C=5P_C-16"

At equilibrium,

"D_A=S_A \\

=>20-2P_A+4P_B+P_C=4P_A-5 \\

=>4P_A+2P_A-4P_B- P_C=20+5"

"=>6P_A-4P_B- P_C=25" ....(1)

"D_B=S_B \\

=> 10+ 3P_A - 5P_B + 2P_C = 3P_B – 7 \\

=>-3P_A+3P_B+5P_B-2P_C=10+7 \\"

"=>-3P_A+8P_B-2P_C=17" ....(2)

"D_C = S_C"

"=>70+4P_A+2P_B-5P_C=5P_C-16 \\

=>-4P_A-2P_B+5P_C+5P_C=70+16"

"=>-4P_A-2P_B+10P_C=86" ....(3)

From the system of equations, we get the determinant as,

"|D|=\begin{vmatrix} 6 &-4 &-1 \\ -3& 8 &-2 \\ -4 &-2 & 10 \end{vmatrix}=6(80-4)+4(-30-8)-1(6+32)=456-152-38=266"

By Cramer's rule,

"P_A^*=\frac{\begin{vmatrix} 25 &-4 &-1 \\ 17& 8 &-2 \\ 86 &-2 & 10 \end{vmatrix}}{|D|}=\frac{25(80-4)+4(170+172)-1(-34-688)}{266} \\

P_A^*=\frac{25 \times 76+4 \times 342-1(-722)}{266}=\frac{1900+1368+722}{266}=\frac{3990}{266} \\

P_A^*=15 \\

Q_A^*=4 \times 15-5=60-5=55 \\

P_B^*=\frac{\begin{vmatrix} 6 &25 &-1 \\ -3& 17 &-2 \\ -4 &86 & 10 \end{vmatrix}}{|D|}=\frac{6(170+172)-25(-30-8)-1(-258+68)}{266} \\

P_B^*=\frac{6 \times 342+25 \times 38-1(-190)}{266}=\frac{2052+950+190}{266}=\frac{3192}{266} \\

P_B^*=12 \\

Q_B^*=3 \times 12-7=36-7=29 \\

P_C^*=\frac{\begin{vmatrix} 6 &-4 &25 \\ -3& 8 &17 \\ -4 &-2 & 86 \end{vmatrix}}{|D|}=\frac{6(688+34)+4(-258+68)+25(6+32)}{266} \\

P_C^*=\frac{6 \times 722-4 \times 190+25 \times 38}{266}=\frac{4332-760+950}{266}=\frac{4522}{266} \\

P_C^*=17 \\

Q_C^*=5 \times 17-16=85-16=69 \\

D_B=10+3P_A-5P_B+2P_C \\

\frac{\mathrm{d} D_B}{\mathrm{d} P_A}=3 \\

\frac{\mathrm{d} D_B}{\mathrm{d} P_B}=-5 \\

\frac{\mathrm{d} D_B}{\mathrm{d} P_C}=2"

At equilibrium,

"D_B=29 , \small P_A^*=15 , \small P_B^*=12 , \small P_C^*=17"

Elasticity of demand with respect to PA,

"\varepsilon _A=\frac{\mathrm{d} D_B}{\mathrm{d} P_A}\frac{P_A}{D_B}=3 \times\frac{15}{29}=1.55"

"\varepsilon _A>0" implies A and B are substitute goods.

Elasticity of demand with respect to PB,

"\varepsilon _B=\frac{\mathrm{d} D_B}{\mathrm{d} P_B}\frac{P_B}{D_B}=-5 \times \frac{12}{29}=-2.0"

"|\varepsilon _B|>1" implies the demand for good B is elastic.

Elasticity of demand with respect to PC,

"\varepsilon _C=\frac{\mathrm{d} D_B}{\mathrm{d} P_C}\frac{P_C}{D_B}=2 \times\frac{17}{29}=1.17"

"\varepsilon _C>0" implies B and C are substitute goods.